The train traveled at 50 MPH from Bedford to Middleton. The trip took 25 minutes. About how far away is Middleton from Bedford?
Answer by Alan3354(69443)

The train traveled at 50 MPH from Bedford to Middleton. The trip took 25 minutes. About how far away is Middleton from Bedford?
-------------
d = r*t
d = 50 mi/hr*25 min*(1hr/60min)
d = 50*25/60 miles
d = 125/6 miles = almost 21 miles